# Incorporating Coupling Knowledge into Echo State Networks for Learning Spatiotemporally Chaotic Dynamics

### v1

@@ Abstract
Machine learning methods have shown promise in learning chaotic dynamical systems, enabling model-free short-term prediction and attractor reconstruction. However, when applied to large-scale, spatiotemporally chaotic systems, purely data-driven machine learning methods often suffer from inefficiencies, as they require a large learning model size and a massive amount of training data to achieve acceptable performance. To address this challenge, we incorporate the spatial coupling structure of the target system as an inductive bias in the network design. Specifically, we introduce physics-guided clustered echo state networks, leveraging the efficiency of the echo state networks as a base model. Experimental results on benchmark chaotic systems demonstrate that our physics-informed method outperforms existing echo state network models in learning the target chaotic systems. Additionally, our models exhibit robustness to noise in training data and remain effective even when prior coupling knowledge is imperfect. This approach has the potential to enhance other machine learning methods.
@@
Chaotic systems are ubiquitous in nature and society. Many such systems, such as those found in the brain and social networks, have governing equations that remain largely unknown. This necessitates data-driven approaches to reconstruct their dynamics from observed time-series data. However, for large-scale systems, purely data-driven methods often require massive amounts of training data and computational resources, highlighting the need to integrate prior knowledge into data-driven methods. In this work, we incorporate prior coupling knowledge from target systems into model design. Our tests suggest that the coupling structure serves as an effective inductive bias for learning spatiotemporally chaotic dynamical systems.
@@ I Introduction
Complex dynamical systems (DSs), which span a wide range of spatial and temporal scales, play a crucial role in various scientific and engineering disciplines, including meteorology, Kashinath et al. (2021) fluid dynamics, Wang et al. (2020) and cosmology Böhmer and Chan (2017). In many cases, such as brain activity and social networks, the governing equations of these systems remain unknown. Understanding the dynamics of chaotic systems is crucial for predicting their behavior, uncovering hidden governing principles, and harnessing their complexity for practical applications in science and technology. These reasons necessitate data-driven approaches to infer their dynamics from observed (ground truth) time-series data. Furthermore, most complex DSs exhibit inherent chaos characterized by the exponential divergence of nearby trajectories,Strogatz (2018) making them challenging to accurately simulate using model-based classical numerical methods with the governing equation. This challenge is further exacerbated in data-driven approaches designed to forecast the system’s future states. Due to the inherent chaoticity in those complex DSs, our focus in the context of DS learning shifts from state forecasting to attractor reconstruction. This entails learning a generative model capable of replicating the system’s invariant (long-term) geometrical and temporal properties.Hess et al. (2023)
@@
Given these challenges, machine learning (ML) methods—particularly recurrent neural networks (RNNs)—have emerged as powerful data-driven approaches for learning chaotic DSs in recent years.Jaeger (2001); Pathak et al. (2018a); Lu, Hunt, and Ott (2018) However, for complex, high-dimensional DSs, purely data-driven learning requires an enormous amount of training data, which is often impractical due to the high cost and difficulty of data collection in many scientific fields. Karniadakis et al. (2021) Additionally, RNN-based methods typically rely on backpropagation through time training, Werbos (1990) which introduces inefficiencies during the unrolling of the computational graph and suffers from vanishing and exploding gradients. Vlachas et al. (2020) Moreover, compared to popular but complex gated RNN models, such as long short-term memory (LSTM) Hochreiter (1998) and gated recurrent unit (GRU), Cho et al. (2014) simpler architectures offer advantages in mathematical tractability and scientific interpretability for learning chaotic systems. Hess et al. (2023) All these reasons highlight the need to integrate prior knowledge into the learning process using a simple ML architecture that balances efficiency and interpretability.Karniadakis et al. (2021); Pathak et al. (2018a, b)
@@
Among the various forms of prior knowledge that can aid DS learning through ML, physical information—such as constraints based on physical models—has gained increasing attention and demonstrated significant potential in recent studies.Wang et al. (2020); Pathak et al. (2018a); Arcomano et al. (2022); Pathak et al. (2018b); Darbon and Meng (2020); Raissi, Perdikaris, and Karniadakis (2019); Sun et al. (2021, 2022a, 2022b) These studies have integrated prior physical knowledge into neural network design by structuring model components accordingly,Wang et al. (2020); Pathak et al. (2018a); Sun et al. (2021) embedding mathematical solutions directly, Pathak et al. (2018b); Arcomano et al. (2022); Darbon and Meng (2020); Sun et al. (2022a, b) or formulating loss functions based on the residual of the governing partial differential equations (PDEs). Raissi, Perdikaris, and Karniadakis (2019) For example, Arcomano et al. Arcomano et al. (2022) proposed a computationally efficient hybrid model that combines a physics-based global atmospheric circulation model with a machine learning component for weather forecasting, achieving improved quantitative forecasting performance.
@@
Building on the above studies, physical knowledge is also expected to contribute to DS learning as a inductive bias for designing specialized structures inside a network layer of an ML model. In this work, we adopt an Echo State Network (ESN) Jaeger (2001) as the base model for DS learning. The ESN, an RNN model within the reservoir computing Verstraeten et al. (2007) paradigm, is well-suited for sequential data processing, particularly in attractor reconstruction, due to its echo state property and efficient training mechanism.Pathak et al. (2017); Jaeger (2001) Inspired by the work of Ref. Pathak et al., 2018a, which introduced a parallelized ESN structure leveraging the local interactions of the target DS in the input layer, we propose a novel model called the physics-guided clustered ESN (PGC-ESN). Our approach leverages the spatial coupling structure of the target DS as prior physical knowledge to design a coupled clustered architecture in the ESN’s reservoir layer, which typically has random connectivity. The effectiveness of multiple clustered structures in ESNs has been studied in Ref. Gallicchio, Micheli, and Pedrelli, 2017; Kawai, Park, and Asada, 2019; Li and Tanaka, 2022. Empirical evaluations demonstrate that the coupling structure inherent in a DS can serve as an inductive bias to guide network design for DS learning. We anticipate that in the future, our work will inspire the development of other artificial neural networks that incorporate prior physical knowledge, not only for DS learning but also for broader ML tasks related to structured data.
@@
The rest of this paper is organized as follows. Section II provides background information on attractor reconstruction, including the definition of a DS and a guide to setting up ESNs to carry out this task. Section III details how we incorporate the coupling structure of a target system into our proposed method. Section IV presents evaluation results for two benchmark spatiotemporal dynamical systems: the Lorenz-96 system,Lorenz (1996) and the Kuramoto–Sivashinsky equation.Kuramoto and Tsuzuki (1976); Ashinsky (1988) The results demonstrate that our proposed method outperforms the standard ESN (Std-ESN), a randomly clustered ESN (RandC-ESN, a non-coupled clustered ESN model we made for comparison), and a parallelized ESN (Paral-ESN) in both long-term attractor reconstruction and short-term prediction when the target system exhibits a sufficient coupling structure. Experiments on the Lorenz-96 system with varied training data noise and increasing nonlinearity further show that our physics-informed methodology enhances the robustness of ESN models to noise and improves their learning performance as the system’s nonlinearity increases. Additionally, we evaluated the PGC-ESN under perturbations to the reservoir’s coupling structure using the Lorenz-96 system and numerically demonstrated that our proposed method remains effective even when the prior coupling knowledge is imperfect. Finally, Section V concludes the paper with a summary of our contributions.
@@ II.1 Dynamical systems
Let $k\geq 1$ be an integer and $U$ be an open subset of $\mathbb{R}^{n}$ representing the system domain. We define the system state $\bm{u}:U\mapsto\mathbb{R}^{m}$ and a variable $\bm{x}\in U$. A differential equation of the form
[…]
The operator $\bm{\mathcal{F}}$ governs the $m$-dimensional DS with respect to the variable $\bm{x}\in\mathbb{R}^{n}$ and can be either linear or non-linear. Nonlinearity is often responsible for the chaotic nature of a DS. Since most DSs evolve over time, and one of the variables in $\bm{x}$ is the time dimension, we commonly use the shorthand notation $\bm{u}(t)$ to represent $\bm{u}(t,x_{1},\dots,x_{n-1})$, where $\{x_{1},\dots,x_{n-1}\}$ are the spatial parts of the variable $\bm{x}$.
Chaotic DSs exhibit a sensitive dependence on initial conditions, meaning that nearby trajectories diverge exponentially over time, making future state forecasting inherently challenging. This exponential divergence rate is called the maximum Lyapunov exponent (MLE), Strogatz (2018) denoted as $\lambda$. Due to their chaotic nature, we focus on reconstructing chaotic DSs in this work, emphasizing the learning of the system’s long-term behavior. For short-term state prediction, we use the Lyapunov time, defined as $T_{\lambda}={\lambda}^{-1}$, as a reference timescale to measure the growth of error between predicted and observed sequences in chaotic systems.
@@ II.2 Dynamical system reconstruction using echo state networks
In the task of attractor reconstruction, the objective is to learn a generative model that approximates the underlying dynamics of an unknown system from observed time-series data. This model should be capable of generating sequences that replicate the long-term behavior of the system, including the geometry of its dynamical structures in state space, such as attractors, and invariant temporal properties, such as the system’s power spectrum, after sufficient training.Hess et al. (2023) In this context, the goal is to implicitly approximate the operator $\mathcal{F}$ on the attractor, which governs the evolution of the underlying DS, as described in Equation 1. To achieve this using an ML model, we sample an observed time series $\bm{u}_{1:T}=\{\bm{u}(1)\dots\bm{u}(T)\}$ from the ground truth system at a fixed step size $\Delta t$, where each $\bm{u}(t)\in\mathbb{R}^{d_{u}}$ is the state at time step $t$ in the state space. The training objective here is a one-step prediction: given the current state $\bm{u}(t)$, the model outputs the prediction $\hat{\bm{u}}(t+1)$ for the next time step.
@@
We provide a schematic illustration of the ESN learning process in Appendix. A, along with the following mathematical explanation for our setup. At time step $t$ of the training phase, the input layer $\hat{\mathbf{R}}_{\mathrm{in}}$ maps the input $\bm{u}(t)$ to the reservoir state $\bm{r}(t)\in\mathbb{R}^{d_{r}}$ linearly by $\hat{\mathbf{R}}_{\mathrm{in}}[\bm{u}(t)]=\bm{W}^{\mathrm{in}}\bm{u}(t)$, where $d_{r}$ is the reservoir size. The readout layer maps the reservoir state to the output $\hat{\bm{u}}(t+1)$ linearly by $\hat{\bm{u}}(t+1)=\hat{\mathbf{R}}_{\mathrm{out}}[\bm{r}(t)]=\bm{W}^{\mathrm{out}}\tilde{\bm{r}}(t)$. The weight matrices $\bm{W}^{\mathrm{in}}\in\mathbb{R}^{{d_{r}}\times{d_{u}}},\bm{W}\in\mathbb{R}^{{d_{r}}\times{d_{r}}}$, and $\bm{W}^{\mathrm{out}}\in\mathbb{R}^{{d_{u}}\times{d_{r}}}$ respectively. Typically, we set the reservoir layer’s dimension ${d_{r}}\gg{d_{u}}$ to ensure that it has sufficient capability to capture the temporal features of the observed state history. The augmented reservoir state $\tilde{\bm{r}}(t)$ is obtained by squaring half of the elements with even indexes in $\bm{r}(t)$ to enrich the reservoir dynamics, which empirically improves the model’s efficiency in attractor reconstruction.Pathak et al. (2018a) The entries of $\bm{W}^{\mathrm{in}}$ are uniformly sampled from $[-\sigma,\sigma]$ and masked randomly to give the matrix a density $P_{\mathrm{in}}$, where $\omega$ is referred to as the input scaling and controls the relative influence of the current input on the reservoir state versus the history of inputs.Lukoševičius (2012) The entries of $\bm{W}$ are uniformly sampled from $[0,1]$, masked randomly to give this matrix a density $P$, and finally rescaled so that its spectral radius (the largest absolute value of the eigenvalues) $\rho<1$. This scaling ensures that the ESN satisfies the echo state property,Jaeger (2001) meaning that the reservoir state $\bm{r}(t)$ becomes independent of its initial conditions as $t$ increases, given a sufficiently long input sequence. The spectral radius $\rho$ also affects the stability of the reservoir activations and the speed at which the influence of past inputs decays in the reservoir.Lukoševičius (2012) We tune $\sigma$ and $\rho$ as hyperparameters and fix the two densities $P_{\mathrm{in}}$ and $P$ in this work.
@@
To learn a DS using an ESN, we can begin by initializing the reservoir state to $\bm{0}\in\mathbb{R}^{d_{r}}$ and warming it up with an observed sequence of length $T_{\mathrm{warm}}$, both in the training and the prediction phases. This process renders the ESN able to encode temporal dependencies on the past state history into the reservoir state $\bm{r}$ and predict multiple sequences without requiring proximity to the training data, highlighting the reusability of the readout layer for generating extended sequences after being trained just once. During the training phase, we feed an observed sequence of length $T_{\mathrm{train}}$ into the model and use the open-loop configuration to evolve the reservoir states as follows:
$\bm{r}(t+1)=f(\bm{W}^{\mathrm{in}}\bm{u}(t)+\bm{W}\bm{r}(t)),$ (2)
@@
where the activation function $f$ is chosen to be the component-wise sigmoid function $f=\tanh$. Note that the open-loop configuration is also used during warm-up. We then record the reservoir states of length $T_{\mathrm{train}}$ and train the readout matrix $\bm{W}^{\mathrm{out}}$ by minimizing the mean squared error (MSE) between the observed and predicted sequences using least-squares regression. To prevent overfitting, a Tikhonov regularization term with parameter $\eta$ is typically added, which penalizes the assignation of large values to the fitting parameters in $\bm{W}^{\mathrm{out}}$. For further regularization, we add Gaussian noise sampled from a normal distribution $\mathcal{N}(0,\kappa\sigma(\bm{u}_{1:T_{\mathrm{Train}}}))$ to the observed training sequences $\bm{u}_{1:T_{\mathrm{Train}}}$, where $\sigma$ denotes the standard deviation of the sequence, and $\kappa$ is the noise level that determines the magnitude of perturbation. To alleviate the RAM requirement and improve the computational efficiency, we use a time-bached schedule, as described in Ref. Vlachas et al., 2020. The loss function for training is as follows:
$\sum_{1\leq t\leq T_{\mathrm{train}}}\lVert\bm{W}^{\mathrm{out}}\mathbf{r}(t)-\bm{u}(t)\rVert^{2}+\eta\lVert\bm{W}^{\mathrm{out}}\rVert^{2}.$ (3)
@@
$\bm{r}(t+1)=f(\bm{W}^{\mathrm{in}}\hat{\bm{u}}(t)+\bm{W}\bm{r}(t)),$ (4)
where $\hat{\bm{u}}(t)=\bm{W}^{\mathrm{out}}\tilde{\bm{r}}(t)$ is the predicted state at time step $t$ generated by the ESN.

### v2

@@ Abstract
Machine learning methods have shown promise in learning chaotic dynamical systems, enabling model-free short-term prediction and attractor reconstruction. However, when applied to large-scale, spatiotemporally chaotic systems, purely data-driven machine learning methods often suffer from inefficiencies, as they require a large learning model size and a massive amount of training data to achieve acceptable performance. To address this challenge, we incorporate the spatial coupling structure of the target system as an inductive bias in the network design. Specifically, we introduce physics-guided clustered echo state networks, leveraging the efficiency of the echo state networks as a base model. Experimental results on benchmark chaotic systems demonstrate that our physics-informed method outperforms existing echo state network models in learning the target chaotic systems. Additionally, we numerically demonstrate that leveraging coupling knowledge into ESN models can enhance their robustness to variations of training and target system conditions. We further show that our proposed model remains effective even when the coupling knowledge is imperfect or extracted directly from time series data. We believe this approach has the potential to enhance other machine-learning methods.
@@
Chaotic systems are ubiquitous in nature and society. Many such systems, such as those found in the brain and social networks, have governing equations that remain largely unknown. This necessitates data-driven approaches to reconstruct their dynamics from observed time-series data. However, for large-scale systems, purely data-driven methods often require massive amounts of training data and computational resources, highlighting the need to integrate system-specific knowledge into data-driven methods. In this work, we incorporate coupling knowledge from target systems into model design. Our tests suggest that the coupling structure serves as an effective inductive bias for learning spatiotemporally chaotic dynamical systems.
@@ I Introduction
Complex dynamical systems (DSs), which span a wide range of spatial and temporal scales, play a crucial role in various scientific and engineering disciplines, including meteorology, Kashinath et al. (2021) fluid dynamics, Wang et al. (2020) and cosmology. Böhmer and Chan (2017) In many cases, such as brain activity and social networks, the governing equations of these systems remain unknown. Understanding the dynamics of chaotic systems is crucial for predicting their behavior, uncovering hidden governing principles, and harnessing their complexity for practical applications in science and technology. These reasons necessitate data-driven approaches to infer their dynamics from observed (ground truth) time-series data. Furthermore, most complex DSs exhibit inherent chaos characterized by the exponential divergence of nearby trajectories, Strogatz (2024) making them challenging to accurately simulate using model-based classical numerical methods with the governing equation. This challenge is further exacerbated in data-driven approaches designed to forecast the system’s future states. Due to the inherent chaoticity in those complex DSs, our focus in the context of DS learning shifts from state forecasting to attractor reconstruction. This entails learning a generative model capable of replicating the system’s invariant (long-term) geometrical and temporal properties. Hess et al. (2023)
@@
Given these challenges, machine learning (ML) methods—particularly recurrent neural networks (RNNs)—have emerged as powerful data-driven approaches for learning chaotic DSs in recent years. Jaeger (2001); Pathak et al. (2018a); Lu, Hunt, and Ott (2018) However, for complex, high-dimensional DSs, purely data-driven learning requires an enormous amount of training data, which is often impractical due to the high cost and difficulty of data collection in many scientific fields. Karniadakis et al. (2021) Additionally, RNN-based methods typically rely on backpropagation through time training, Werbos (1990) which introduces inefficiencies during the unrolling of the computational graph and suffers from vanishing and exploding gradients. Vlachas et al. (2020) Moreover, compared to popular but complex gated RNN models, such as long short-term memory (LSTM) Hochreiter (1998) and gated recurrent unit (GRU), Cho et al. (2014) simpler architectures offer advantages in mathematical tractability and scientific interpretability for learning chaotic systems. Hess et al. (2023) All these reasons highlight the need to integrate system-specific knowledge into the learning process using a simple ML architecture that balances efficiency and interpretability. Karniadakis et al. (2021); Pathak et al. (2018a, b)
@@
Among the various forms of system-specific knowledge that can aid DS learning through ML, physical information—such as constraints based on physical models—has gained increasing attention and demonstrated significant potential in recent studies. Wang et al. (2020); Pathak et al. (2018a); Arcomano et al. (2022); Pathak et al. (2018b); Darbon and Meng (2020); Raissi, Perdikaris, and Karniadakis (2019); Sun et al. (2021, 2022a, 2022b) These studies have integrated prior physical knowledge into neural network design by structuring model components accordingly, Wang et al. (2020); Pathak et al. (2018a); Sun et al. (2021) embedding mathematical solutions directly, Pathak et al. (2018b); Arcomano et al. (2022); Darbon and Meng (2020); Sun et al. (2022a, b) or formulating loss functions based on the residual of the governing partial differential equations (PDEs). Raissi, Perdikaris, and Karniadakis (2019) For example, Arcomano et al. Arcomano et al. (2022) proposed a computationally efficient hybrid model that combines a physics-based global atmospheric circulation model with a machine learning component for weather forecasting, achieving improved quantitative forecasting performance.
@@
Building on the above studies, physical knowledge is also expected to contribute to DS learning as a inductive bias for designing specialized structures inside a network layer of an ML model. In this work, we adopt an Echo State Network (ESN) Jaeger (2001) as the base model for DS learning. The ESN, an RNN model within the reservoir computing Verstraeten et al. (2007) paradigm, is well-suited for sequential data processing, particularly in attractor reconstruction, due to its echo state property and efficient training mechanism. Pathak et al. (2017); Jaeger (2001) Inspired by the work of Ref. Pathak et al., 2018a, which introduced a parallelized ESN structure leveraging the local interactions of the target DS in the input layer, we propose a novel model called the physics-guided clustered ESN (PGC-ESN). Our approach leverages the spatial coupling structure of the target DS as prior physical knowledge to design a coupled clustered architecture in the ESN’s reservoir layer, which typically has random connectivity. The effectiveness of multiple clustered structures in ESNs has been studied in Ref. Gallicchio, Micheli, and Pedrelli, 2017; Kawai, Park, and Asada, 2019; Li and Tanaka, 2022.
In practical scenarios, prior knowledge of the target system’s coupling structure is typically imperfect or even unavailable. To simulate the impact of using guidance from imperfect prior knowledge, we evaluate PGC-ESN with a perturbed coupling structure. Inspired by the work of Ref. Goldmann et al., 2022, which extracted the delay of the Mackey–Glass system from time series data to enhance the predictive ability of a model known as delayed echo state network, we also extract the coupling structure directly from time series data and use this information to guide the design of PGC-ESN.
@@
Empirical evaluations on two benchmark spatiotemporal dynamical systems, the Lorenz-96 system Lorenz (1996) and the Kuramoto–Sivashinsky equation Kuramoto and Tsuzuki (1976); Ashinsky (1988), demonstrate that when the target system exhibits a sufficient coupling structure, our proposed method outperforms the standard ESN (Std-ESN), a randomly clustered ESN (RandC-ESN, a non-coupled clustered ESN model we made for comparison), and a parallelized ESN (Paral-ESN) in long-term attractor reconstruction under limited model capacity, and in short-term prediction even when the model size is sufficiently large. Experiments on the Lorenz-96 system with varying training noise level, training data size, sequence temporal resolution, and increasing system nonlinearity further show that the physics-informed methodologies used in PGC-ESN and Paral-ESN can enhance the robustness of ESN models to varied learning settings and improve their learning performance even when the target system’s nonlinearity increases. Additionally, we demonstrate the effectiveness of PGC-ESN numerically on the Lorenz-96 system when provided with imperfect prior or extracted coupling knowledge, highlighting the practical applicability of our method in scenarios where prior knowledge of the target system is imperfect or even unavailable.
Overall, we numerically show that the coupling structure inherent in a DS can serve as an inductive bias to guide network design for DS learning. We anticipate that in the future, our work will inspire the development of other artificial neural networks that incorporate system-specific knowledge, not only for DS learning but also for broader ML tasks related to structured data.
The rest of this paper is organized as follows. Section II provides background information on attractor reconstruction, including the definition of a DS and a guide to setting up ESNs to carry out this task. Section III details how we incorporate the coupling structure of a target system as prior or extracted physical knowledge into our proposed method. Section IV presents the evaluation results. Finally, Section V concludes the paper with a summary of our contributions and several promising directions for future research.
@@ II.1 Dynamical systems
Let $k\geq 1$ be an integer and $U$ be an open subset of $\mathbb{R}^{n}$ representing the system domain. We define the system state $\bm{u}:U\rightarrow\mathbb{R}^{m}$ and a variable $\bm{x}\in U$. An expression of the form
$\mathcal{F}(D^{k}\bm{u}(\bm{x}),D^{k-1}\bm{u}(\bm{x}),\dots,D\bm{u}(\bm{x}),\bm{u}(\bm{x}),\bm{x})=\bm{0}$ (1)
is called a $k^{\text{th}}$-order system of partial (or ordinary when $n=1$) differential equations, where $\mathcal{F}:\mathbb{R}^{mn^{k}}\times\mathbb{R}^{mn^{k-1}}\times\dots\times\mathbb{R}^{mn}\times\mathbb{R}^{m}\times U\rightarrow\mathbb{R}^{m}$, and $D$ denotes either a partial or ordinary derivative operator.
The operator $\mathcal{F}$ governs the $m$-dimensional DS with respect to the variable $\bm{x}\in\mathbb{R}^{n}$ and can be either linear or non-linear. Nonlinearity is often responsible for the chaotic nature of a DS. Since most DSs evolve over time, and one of the variables in $\bm{x}$ is the time dimension, we commonly use the shorthand notation $\bm{u}(t)$ to represent $\bm{u}(t,x_{1},\dots,x_{n-1})$, where $\{x_{1},\dots,x_{n-1}\}$ are the spatial parts of the variable $\bm{x}$.
Chaotic DSs exhibit a sensitive dependence on initial conditions, meaning that nearby trajectories diverge exponentially over time, making future state forecasting inherently challenging. This exponential divergence rate is called the maximum Lyapunov exponent (MLE), Strogatz (2024) denoted as $\lambda$. Due to their chaotic nature, we focus on reconstructing chaotic DSs in this work, emphasizing the learning of the system’s long-term behavior. For short-term state prediction, we use the Lyapunov time, defined as $T_{\lambda}={\lambda}^{-1}$, as a reference timescale to measure the growth of error between predicted and observed sequences.
@@ II.2 Dynamical system reconstruction using echo state networks
In the task of attractor reconstruction, the objective is to learn a generative model that approximates the underlying dynamics of an unknown system from observed time-series data. This model should be capable of generating sequences that replicate the long-term behavior of the system, including the geometry of its dynamical structures in state space, such as attractors, and invariant temporal properties, such as the system’s power spectrum, after sufficient training. Hess et al. (2023) In this context, the goal is to implicitly approximate the operator $\mathcal{F}$ on the attractor, which governs the evolution of the underlying DS, as described in Equation 1. To achieve this using an ML model, we sample an observed time series $\bm{u}_{1:T}=\{\bm{u}(1)\dots\bm{u}(T)\}$ of length $T$ from the ground truth system at a fixed integration step size $\Delta t$, where each $\bm{u}(t)\in\mathbb{R}^{d_{u}}$ is the state at time step $t$ in the state space. The training objective here is a one-step prediction: given the current state $\bm{u}(t)$, the model outputs the prediction $\hat{\bm{u}}(t+1)$ for the next time step.
@@
We provide a schematic illustration of the ESN learning process in Appendix A.1, along with the following mathematical explanation for our setup. At time step $t$ of the training phase, the input layer $\hat{\mathbf{R}}_{\mathrm{in}}$ embeds the input $\bm{u}(t)$ into the reservoir state $\bm{r}(t)\in\mathbb{R}^{d_{r}}$ linearly by $\hat{\mathbf{R}}_{\mathrm{in}}[\bm{u}(t)]=\bm{W}^{\mathrm{in}}\bm{u}(t)$, where $d_{r}$ is the reservoir size. The readout layer maps the reservoir state to the output $\hat{\bm{u}}(t+1)$ linearly by $\hat{\bm{u}}(t+1)=\hat{\mathbf{R}}_{\mathrm{out}}[\bm{r}(t)]=\bm{W}^{\mathrm{out}}\tilde{\bm{r}}(t)$. The weight matrices $\bm{W}^{\mathrm{in}}\in\mathbb{R}^{{d_{r}}\times{d_{u}}},\bm{W}\in\mathbb{R}^{{d_{r}}\times{d_{r}}}$, and $\bm{W}^{\mathrm{out}}\in\mathbb{R}^{{d_{u}}\times{d_{r}}}$ respectively. Typically, we set the reservoir layer’s dimension ${d_{r}}\gg{d_{u}}$ to ensure that it has sufficient capability to capture the temporal features of the observed state history. The augmented reservoir state $\tilde{\bm{r}}(t)$ is obtained by squaring half of the elements with even indexes in $\bm{r}(t)$ to enrich the reservoir dynamics, which empirically improves the model’s efficiency in attractor reconstruction. Pathak et al. (2018a) The entries of $\bm{W}^{\mathrm{in}}$ are uniformly sampled from $\beta,\beta]$ and masked randomly to give the matrix a density $p_{\mathrm{in}}$, where $\beta$ is referred to as the input scaling and controls the relative influence of the current input on the reservoir state versus the history of inputs. Lukoševičius (2012) The entries of $\bm{W}$ are uniformly sampled from $[0,1]$, masked randomly to give this matrix a density $p$, and finally rescaled so that its spectral radius (the largest absolute value of the eigenvalues) $\rho<1$. This scaling ensures that the ESN satisfies the echo state propertyJaeger (2001) meaning that the reservoir state $\bm{r}(t)$ becomes independent of its initial conditions as $t$ increases, given a sufficiently long input sequence. The spectral radius $\rho$ also affects the stability of the reservoir activations and the speed at which the influence of past inputs decays in the reservoir. Lukoševičius (2012) In our experiments, we tune the spectral radius $\rho$ and the input scaling $\beta$ as hyperparameters. The two densities, $p_{\mathrm{in}}$ and $p$, are fixed when prior coupling knowledge of the target system is available, and are tuned otherwise.
@@
To learn a DS using an ESN, we can begin by initializing the reservoir state to $\bm{0}\in\mathbb{R}^{d_{r}}$ and warming it up with an observed sequence of length $T_{\mathrm{warm}}$, both in the training and the prediction phases. This process renders the ESN able to encode temporal dependencies on the past state history into the reservoir state $\bm{r}$ and predict multiple sequences without requiring proximity to the training data, highlighting the reusability of the readout layer for generating extended sequences after being trained just once. Vlachas et al. (2020) During the training phase, we feed an observed sequence of length $T_{\mathrm{train}}$ into the model and use the open-loop configuration to evolve the reservoir states as follows:
$\bm{r}(t)=f(\bm{W}^{\mathrm{in}}\bm{u}(t)+\bm{W}\bm{r}(t-1)),$ (2)
@@
where the activation function $f$ is chosen to be the component-wise sigmoid function $f=\tanh$. Note that the open-loop configuration is also used during warm-up. We then record the reservoir states of length $T_{\mathrm{train}}$ and train the readout matrix $\bm{W}^{\mathrm{out}}$ by minimizing the mean squared error (MSE) between the observed and predicted sequences using least-squares regression. To prevent overfitting, a Tikhonov regularization term with parameter $\eta$ is typically added, which penalizes the assignation of large values to the fitting parameters in $\bm{W}^{\mathrm{out}}$. For further regularization, we add Gaussian noise sampled from a normal distribution $\mathcal{N}\left(\mathbf{0},\,\kappa^{2}\operatorname{diag}\left(\bm{\sigma}(\bm{u}_{1:T_{\mathrm{Warm}}+T_{\mathrm{Train}}})^{2}\right)\right)$ to the observed training sequences $\bm{u}_{1:T_{\mathrm{Warm}}+T_{\mathrm{Train}}}$, where $\bm{\sigma}$ denotes the vector of standard deviations over time, computed per state-space dimension of the sequence, and $\kappa$ is the noise level that determines the magnitude of perturbation. To alleviate the RAM requirement and improve the computational efficiency, we use a time-batched schedule, as described in Ref. Vlachas et al., 2020. The loss function for training is as follows:
$\sum_{1\leq t\leq T_{\mathrm{train}}}\lVert\hat{\bm{u}}(t)-\bm{u}(t)\rVert^{2}+\eta\lVert\bm{W}^{\mathrm{out}}\rVert^{2}.$ (3)
@@
$\bm{r}(t)=f(\bm{W}^{\mathrm{in}}\hat{\bm{u}}(t)+\bm{W}\bm{r}(t-1)),$ (4)
where $\hat{\bm{u}}(t)=\bm{W}^{\mathrm{out}}\tilde{\bm{r}}(t-1)$ is the predicted state at time step $t$, generated by the ESN using the trained $\bm{W}^{\mathrm{out}}$.